# DiffHO: Diffusion-Based Synthesis of Hand-Object Interactions from Textual Descriptions

### v1

@@ 3. Denoising Diffusion Probabilistic Models
Conditional diffusion models generate output conditioned on various kinds of input such as text (Rombach et al., 2022) and other semantic information (Zhang et al., 2023). Such a model is obtained by providing the condition as additional input to the auto-encoder resulting in outputs , or , where c is the conditional input encoding.
3.1. Guiding the diffusion model
@@
Diffusion models can be guided to generate outputs of the required form. Two popular methods to guide the diffusion models are classifier guidance (Dhariwal and Nichol, 2021) and classifier-free guidance (Ho and Salimans, 2022). We describe here the classifier-guidance method which is more relevant to our work and refer the reader to (Ho and Salimans, 2022) for details on classifier-free guidance.
The classifier-guidance method approximates , where c is the guidance, by a Gaussian:
@@
Intuitively, the scaled gradient of the guidance signal, is used to nudge the denoised mean to generate samples in accordance with the guidance c. could represent the class probability which can be implemented using a neural-net (Dhariwal and Nichol, 2021), or could represent a cost term that is optimized (Karunratanakul et al., 2023). The classifier guidance method has the advantage that any guidance signal that is differentiable can be used to guide the diffusion and can use pre-trained diffusion models without retraining them.
@@ 4. Method
We focus on the task of generating a sequence of hand and object poses , where indicate a sequence of left and right hand poses, and is a sequence of object poses given some input conditioning signal c, which in our case comprises a text description and the object mesh. In other words, we want to model a conditional probability distribution with a motion DDPM. The optional term is a differentiable goal function that should be minimized, e.g., the distance to a target grasp. See Fig. 2 for an overview.
In Sec. 4.1, we introduce a compact representation that tightly couples the hand and object poses and helps in generating realistic hand-object interactions. We propose a two-stage generation of hand-object interactions in Sec. 4.2 and grasp guidance in Sec. 4.3. Lastly, in Sec. 4.4, we introduce our newly-collected detailed textual annotations for the GRAB dataset that enable more controlled generation of hand-object interactions.
@@ 4.1. Canonicalized Hand-Object Representation
Our proposed representation of length N contains information about both hands and an object. In particular, the object poses are defined as , where is a sequence of 3D object locations, with , and is a sequence of 3D object orientations, with given in 6D representation (Zhou et al., 2019). The hands are represented by: , where denotes the left and denotes the right hand. This representation is based on the parametric MANO hand model (Romero et al., 2017). Specifically, the global 3D positions of the hands are given by and the global 3D orientations are represented by . To achieve more natural poses, we define the local hand pose in a PCA space of the MANO model with the first 24 components as . Lastly, denotes distances between each hand joint (including the fingertips) and its closest point on the object mesh. As we show in our ablations (see Tab. 4) this representation reduces physical artifacts such as interpenetration. We first normalize the object positions to be at the origin and then adjust the hands’ positions accordingly. Hence, all position components are relative to the object origin at the initial frame. Due to the limited data, such a representation performs better than a per-frame object-relative representation (Christen et al., 2022; Zhang et al., 2024) as we show in 5.6.
@@ 4.2. Two-Stage Hand-Object Generation
We propose to split the synthesis into grasping and interaction stages to facilitate generalization to unseen objects. The grasping stage models the motion from an initial pose to an object grasp. The interaction stage performs an action-based manipulation of the object. Our two-phase approach is motivated by the fact that any hand-object interaction involves an approaching phase where one or two hands grasp a static object, followed by an intent-driven interaction with the object. By decoupling the action-based manipulation from grasping, we can leverage the entire set of motions in the dataset for training the grasping phase, irrespective of the action that is performed in the interaction phase. This insight helps counteract the limited scale of the dataset.
As shown in Fig. 2, our grasping and interaction diffusion models are conditioned on three types of inputs to encode intent, namely, 1) text prompts describing the object and the action to be carried out, 2) object mesh which provides the geometry of the object, and 3) the time step in the denoising process. We use CLIP (Radford et al., 2021) embeddings to encode the text prompt denoted by , BPS (Prokudin et al., 2019) to encode the object mesh denoted by and an MLP to encode the time step as in previous works (Karunratanakul et al., 2023; Dhariwal and Nichol, 2021). We provide more implementation details in supplementary material and describe the grasping and interaction phases below in detail.
@@ 4.2.1. Grasping Stage
The grasping phase is defined as a sequence containing one or two hands approaching a static object from a rest pose until the object is grasped. As the object is static throughout this phase we do not predict the object motion and only model the hand pose sequence. In order to obtain the training data for this phase we use heuristics based on object poses to determine sequence boundaries from the larger action sequences and normalize the sequence lengths by interpolating or downsampling the poses. The text description in this phase is a generic sentence (”The person grasps the ¡object¿.”).
We use the diffusion model that directly predicts the noise (see Section 3) in classifier-free manner (Ho and Salimans, 2022) which is more susceptible to inference-time guidance as shown in GMD (Karunratanakul et al., 2023). In other words, our grasping diffusion model training loss is given by
where indicates the synthesized motion at diffusion step .
@@ 4.2.2. Interaction Stage
This phase comprises the motion of both hands and the object after a grasp has been established. In this phase, the hands manipulate and interact with the object according to the action defined through the textual prompt. We obtain the training data for this phase in a similar way as the grasping phase based on object pose heuristics. We employ a diffusion model in a classifier-free manner (see Section 3), i.e., we directly predict the denoised output. This is shown to result in less noisy, higher quality motion than (but less susceptibility to guidance) (Karunratanakul et al., 2023), which is critical during the interaction phase where the hand-object interaction takes place. Our training loss is given by
@@ 4.2.3. Grasp to Interaction Stage Transition
Maintaining a seamless transition between the grasping and interaction outputs is critical to generating realistic motion sequences. We rely on inference-time guidance (Karunratanakul et al., 2023) and motion imputing on the interaction model to achieve a smooth transition. To this end, the interaction diffusion model is trained to generate the complete sequence (including grasping). We then impute the entire sequence generated by the (guided) grasping model to the start of the interaction sequence and apply guidance. To perform the imputation we define respectively the projections and that resize the grasp reference or sub-sequence poses to that of by filling in zeros. Similarly, and represents the imputation regions of the grasp reference and sub-sequence on . At each denoising step, the imputated denoised output, , is given by
where and represents elementwise multiplication. We use the classifier guidance mechanism described in Sec. 3.1 and approximate the gradient signal for sub-sequence guidance as
where denotes the set of grasping poses and denotes the generated grasping poses from the grasping diffusion model. We show in Sec. 5.6 that this technique leads to a smoother transition between pre- and interaction compared to a single frame transition.
@@ 4.3. Grasp Guidance
To add controllability to the generation and improve generalization to unseen objects, we introduce grasp guidance for the grasping phase, which can be optionally be added at inference time. We leverage a single grasp frame, defined as the last frame of the grasping phase, that provides a prior about how and with which hand an object should roughly be grasped. Such a grasp can be obtained from motion capture or from an off-the-shelf hand-object pose estimator (Pavlakos et al., 2023) as demonstrated in Sec. 5.8. We provide inference-time guidance to our trained diffusion model to grasp the object in a pre-defined manner. More specifically, we approximate the gradient of the guidance signal in Eq. 4 as
where is the hand(s) pose in the last frame of the grasping sequence, and is the pre-defined grasp pose.
| Method | Backbone | SD [] () | OD []() | IV [] () | ID [] () | CR | MM () | DIV () |
Real Mocap (GRAB) | - | - | 0.183 | 5.65 | 5.89 | 0.08 | 0.22 | 1.06 | 0.708 | 0.729 | ||||
| IMoS (Ghosh et al., 2023) | CVAE | 0.002 | 0.149 | 7.14 | 11.47 | 0.05 | 0.25 | 1.02 | 0.579 | 0.588 | |||||
| DiffHO | Transformer | 0.088 | 0.185 | 6.65 | 8.39 | 0.067 | 0.33 | 1.09 | 0.760 | 0.810 | |||||
| DiffHO | UNet | 0.109 | 0.188 | 6.02 | 7.92 | 0.064 | 0.30 | 1.08 | 0.833 | 0.875 | |||||
Real Mocap (GRAB) | - | - | 0.183 | 3.40 | 5.63 | 0.071 | 0.20 | 1.08 | 0.683 | 0.784 | ||||
| IMoS* (Ghosh et al., 2023) | CVAE | 0.002 | 0.132 | 10.38 | 12.45 | 0.048 | 0.22 | 1.05 | 0.561 | 0.581 | |||||
| DiffHO | Transformer | 0.133 | 0.185 | 7.99 | 10.87 | 0.073 | 0.24 | 1.09 | 0.750 | 0.803 | |||||
| DiffHO | UNet | 0.134 | 0.179 | 9.03 | 11.39 | 0.086 | 0.23 | 1.09 | 0.755 | 0.837 |
4.4. Textual Augmentations
@@
Our model generates hand-object interactions based on textual inputs. To the best of our knowledge, there is currently no hand-object interaction dataset that contains detailed textual descriptions. For instance, the GRAB dataset (Taheri et al., 2020) only provides categorical action labels. To address this limitation, we auto-generated sentences using the template “the person ¡verb¿ + ¡object¿” (Ghosh et al., 2023) (referred to as “simple” text). However, these sentences lack detailed information, which limits the controllability of the model via textual inputs. Therefore, we contribute carefully annotated textual descriptions of the GRAB dataset (referred to as “detailed” sentences). We instructed our annotators to carefully watch each ground truth video and describe the actions in three distinct stages of pre-action, action, and post-action. These descriptions are further augmented with hand and positional information. An example from the dataset is as follows: ”The person picks up the apple with the right hand, passes it to their left hand, and then places it back with their right hand”. These detailed descriptions enhance the dataset, enabling more accurate evaluation and precise control of hand-object interactions as we show in Sec. 5.7.
@@ 5.1. Baselines
Our method focuses on synthesizing detailed hand-object motion based on text input. To the best of our knowledge, DiffHO is the first method to tackle this problem. While there is no direct baseline for hand-object motion synthesis based on language, the closest to our approach is IMoS (Ghosh et al., 2023), which focuses on text-based whole-body human-object interaction synthesis. IMoS first generates a body articulation based on an instruction label (action + object), and then optimizes for the object pose using a grasp heuristic. Since our paper focuses on hand articulations and object motions, we omit the full-body movements from our evaluations.
To compare our method against human-body diffusion models, we adopted these methods to the HOI setting. All variants use our proposed representation. MDM (Tevet et al., 2023) does not use guidance, and for GMD (Karunratanakul et al., 2023) we use gradients to guide the model towards the grasp frame, but omit the 2D-trajectory generation stage because it is specific to human motions.
@@ 5.2. Data
GRAB: We utilize the subject-based split of the GRAB dataset (Taheri et al., 2020) proposed in IMoS to run a direct comparison. However, as this split does not contain unseen objects, we also create a new split featuring unseen objects based on the criteria of object similarity and semantics (see supplemental material for details).
HO-3D: To further evaluate guidance and object generalization, we estimate hand-object poses on the HO-3D dataset (Hampali et al., 2020) and guide our model to an estimated target pose.
@@ 5.3. Evaluation Metrics
Following previous works on motion synthesis (Ghosh et al., 2023; Tendulkar et al., 2023; Taheri et al., 2022; Braun et al., 2024), we report several physics-based metrics and evaluate our model’s accuracy through metrics that assess motion diversity and action features.
@@
Physics Based Metrics: We report the interpenetration volume (IV) as the amount of MANO vertices that penetrate the object mesh and the maximum interpenetration depth (ID). Furthermore, we compute the contact ratio (CR) as the average ratio of hand vertices that are in contact, which is defined as points within a 5mm threshold of the object mesh.
@@
Motion Diversity Metrics: We report the sample diversity (SD) between wrist trajectories in motion space. We sample the same input conditions five times and compute the pairwise Euclidean distance between the samples. We then report the mean sample diversity over all test prompts. We also provide the overall diversity (OD), which measures the pairwise Euclidean distance between all test samples. Since the wrist positions are normalized with respect to the initial object pose, no further normalization is applied.
@@
Guidance Metrics: To assess how well the models adhere to the guidance signal, we report the mean error between the predictions and the reference grasp (GE), the ratio at which the diffusion model maintains the correct handedness according to the reference (HA). Furthermore, we evaluate the wrist velocities at the transition between grasping and interaction phase ().
Action Feature Based Metrics: To assess the naturalness and variety of the generated motions, we use the action recognition accuracy, diversity, and multimodality metrics, following (Ghosh et al., 2023). We provide additional metrics (e.g. FID, KID), standard deviations and details about how these metrics are computed in our supplemental material.
5.4. Comparison to the State-of-the-Art
We conduct a set of experiments to compare to relevant baselines quantitatively. To demonstrate the generality of our method, we present qualitative examples of our method in Fig. 3 and supplementary materials (along with failure cases).
Setting 1 - Interaction Only: We compare our approach against the state-of-the-art human-object motion synthesis approach of IMoS in their postgrasp-only setting and report results in Tab. 1. To ensure a fair comparison in this experiment, no grasping and neither grasp guidance nor our detailed text annotations are used. We test our hand-object representation with two backbones, MDM’s transformer (Tevet et al., 2023) and GMD’s UNet (Karunratanakul et al., 2023). Our model significantly outperforms IMoS across all physics and motion diversity metrics as well as on the action recognition accuracies. In particular, our method yields motions that are significantly more diverse (higher SD and OD), exhibit less interpenetration (lower IV and ID), and encode motion features that align with action type (higher action recognition accuracies). Note that IMoS assumes the availability of a ground-truth test motion to initialize the first few frames with it, while ours predicts the entire sequence. Yet, our approach achieves considerable performance improvements. While IMoS provides compelling realistic full body synthesis results, fine-grained finger motions as well as post-optimized object motions lag behind the realism and diversity demonstrated by our method, as shown in Fig. 3 and in our supplementary video. The higher scores of DiffHO compared to real mocap from the GRAB dataset in Tab. 1 can occur because the action classifier is trained on GRAB’s training-split, which is the distribution our model covers,whereas some of the test sequences in GRAB (real mocap) have quite distinct motions that the classifier may misclassify. However, the high scores demonstrate the representative power of DiffHO for generating realistic hand-object sequences that capture inherent motion and action semantics. Our method with a UNet backbone outperforms a transformer backbone in most metrics, which is in line with GMD’s findings that UNet works better in low-data regimes. To assess the action recognition accuracy, we train two models, one that uses only hand 3D joints as input and another incorporating object 3D position as additional input. We demonstrate that DiffHO’s synthesized object motions contribute to larger performance improvements compared to the post-optimized motions computed by IMoS ( improvement for ours vs for IMoS) as shown in Tab. 1. This confirms the benefits of jointly modeling hand-object motion instead of decoupling it into separate components.
@@ 5.4. Comparison to the State-of-the-Art
Setting 2 - Grasping + Interaction: In this experiment, we switch to the setting of generating the whole sequence of grasping and interaction, and are interested how well the models can be guided while generating high quality motions (without physical artifacts). We report results in Tab. 2 and find that our method outperforms baselines from human motion diffusion. In particular, our method better adheres to the grasp frame in terms of grasp error (GE) and handedness accuracy (HA), and achieves more natural grasps (lower IV). This justifies our contributions of two-stage separation and subsequence guidance over previous diffusion models.
5.5. Perceptual Study
To evaluate the visual quality of our motions, we conducted a perceptual user study on a set of 21 participants, who were randomly assigned one of four sets, each consisting of 30 randomly selected sequences. We compared our approach against IMoS. We displayed results from our method and IMoS side-by-side in random order, along with the input sentence description used to synthesize motions. We asked the following questions to the participants: “Which sequence is more realistic?” and “Which sequence has more more variety in motion?”. We defined a sequence to be more realistic if the overall motion looks human-like, e.g., if the objects are grasped realistically and there are less artifacts such as floating objects or interpenetration. Moreover, a sequence has more variety if the object is manipulated multiple times and in distinct ways. In 63.1 of the responses, participants selected our method as the most realistic compared to IMoS. Even more distinctively, in 72.9 of the responses, participants favored our method to contain more variety than IMoS. We obtained statistically significant Fleiss’ kappa scores (Fleiss, 1971) of 0.34 and 0.43 for Realism and Diversity, respectively, at p¡0.05. Please see supplemental material for more details.

### v2

@@ 3. Preliminary: Diffusion Models
where is either , , or . Depending on the application, a different choice of the model output works better in practice, e.g., directly predicting the output for human motion (Tevet et al., 2023) or noise prediction for images (Rombach et al., 2022) or guidance (Karunratanakul et al., 2023). Conditional diffusion models generate output conditioned on various kinds of input such as text (Rombach et al., 2022) or other semantic information (Zhang et al., 2023). Such a model is obtained by providing the condition as additional signal, c, to the auto-encoder, resulting in outputs , , or .
@@
Diffusion models can be guided to generate outputs of the required form. Two popular methods to guide the diffusion models are classifier guidance (Dhariwal and Nichol, 2021) and classifier-free guidance (Ho and Salimans, 2022). We describe the classifier-guidance method here and refer the reader to (Ho and Salimans, 2022) for details on classifier-free guidance. The classifier-guidance method approximates , where c is the guidance signal. Deriving from this, Dhariwal and Nichol (2021) show that the new sampling process can be expressed as
@@
where indicates the original mean, is the scaling of the gradients, and is a variance scheduler. Here the mean is shifted by the scaled gradient. Intuitively, the scaled gradient of the guidance signal, nudges the denoised mean to generate samples in accordance with the guidance c. could represent the class probability which can be implemented using a neural-net, or could represent a cost term that is optimized. The classifier guidance method has the advantage that it can use pre-trained diffusion models without retraining them and can benefit from any guidance signal that is differentiable to guide the diffusion.
@@ 4. Method
We focus on the task of generating a sequence of hand and object poses , where indicate a sequence of left and right hand poses, and is a sequence of object poses given some input conditioning signal c, which in our case comprises a text description and the object mesh. In other words, we want to model a conditional probability distribution with a motion DDPM. The optional term can be any differentiable goal function that should be minimized, e.g., the distance to a target grasp in our case (see Sec. 4.3). See Fig. 2 for an overview. In Sec. 4.1, we introduce our compact representation that couples the hand and object poses. We propose a two-stage generation of hand-object interactions in Sec. 4.2 and grasp guidance in Sec. 4.3. Lastly, in Sec. 4.4, we introduce our newly-collected detailed textual annotations for GRAB that enable more controlled generation of HOI.
@@ 4.1. Canonicalized Hand-Object Representation
Our proposed canonicalized representation , where is the number of steps in the sequence, contains information about both hands and an object. The object poses are defined as , where is a sequence of 3D object locations, with , and is a sequence of 3D object orientations, with in 6D representation (Zhou et al., 2019). The hands are represented by: , where denotes the left and the right hand, respectively. This representation is based on the parametric MANO hand model (Romero et al., 2017). Specifically, the global 3D positions of the hands are given by and the global 3D orientations are represented by . To achieve more natural poses, we define the local hand pose in PCA space of MANO with the first 24 components as . We further encode signed distances, , between each hand joint and its closest point on the object mesh. As we show in ablations (see Tab. 4), this approximation of a signed-distance field (SDF), coupled with our diffusion model reduces physical artifacts. Existing works (Christen et al., 2022) encode hand-object poses with a per-frame object-relative representation, which tightly couples the hands to the training object pose and therefore limits generalization to unseen objects. In contrast, we represent hand positions relative to the normalized object position at the frame that marks the boundary between the grasp and interaction stage. This representation is less variant to specific training object pose and more correlated with the global poses involved in HOI, which is often shared across different training objects in HOI datasets (Chao et al., 2021; Taheri et al., 2020). As a result, it yields better generalization and accuracy for HOI across different (unseen) objects, as we show in ablations in Sec. 5.6.
@@ 4.2. Two-Stage Hand-Object Generation
We split the synthesis into grasping and interaction stages to facilitate generalization to unseen objects. The grasping stage models the motion from an initial pose to object grasp. The interaction stage performs an action-based manipulation of the object. Our two-phase approach is motivated by the fact that any hand-object interaction involves an approaching phase where one or two hands grasp a static object, followed by an intent-driven interaction with the object. By decoupling the action-based manipulation from grasping, we can leverage the entire set of motions in the data for training the grasping phase, irrespective of the action that is performed in the interaction phase. These insights help counteract the limited scale of the data and improve generalization to different textual inputs.
As shown in Fig. 2, our grasping and interaction diffusion models are conditioned on three types of inputs, namely, 1) text prompts describing the object and the action to be carried out, 2) object mesh which provides the geometry of the object, and 3) the time step in the denoising process. We use CLIP (Radford et al., 2021) embeddings to encode the text prompt denoted by , BPS (Prokudin et al., 2019) to encode the object mesh denoted by and an MLP to encode the time step as in previous works (Dhariwal and Nichol, 2021). We provide more implementation details in supplementary material and describe the grasping and interaction phases below in detail.
@@ 4.2.1. Grasping Stage
This stage is defined as a sequence containing one or two hands approaching a static object from a rest pose until the object is grasped. As the object is static throughout this phase we do not predict the object motion and only model the hand pose sequence. To obtain the training data for this phase we use heuristics to determine sequence boundaries from the larger action sequences. In particular, the boundary is defined as the first frame where the object’s lateral and vertical velocity surpass a threshold of 0.01m/s and at least 7 vertices of the hand are in contact with the object . We then normalize the sequence lengths by interpolating or downsampling the poses. The text description in this phase is a generic sentence (”The person grasps the ¡object¿.”).
We use the diffusion model that directly predicts the noise in classifier-free manner (Ho and Salimans, 2022) which is more suitable for inference-time guidance (Karunratanakul et al., 2023), e.g., grasp guidance (Section 4.2.3). In other words, our grasping diffusion model training loss is given by
where is the synthesized motion at diffusion step and the final output of the grasping stage after denoising with is .
@@ 4.2.2. Interaction Stage
This stage comprises the motion of both hands and the object after a grasp has been established. The hands manipulate and interact with the object according to the action defined through the textual prompt. We use the same sequence boundary determined in the grasping stage to get the training data for interaction. We employ a diffusion model in a classifier-free manner (see Section 3), i.e., we directly predict the denoised output , that result in less noisy, higher quality motion than diffusion model for the interaction phase that does not involve inference-time grasp guidance. Our training loss is given by
@@ 4.2.3. Grasp to Interaction Stage Transition
Maintaining a seamless transition between the grasping and interaction outputs is critical to generating realistic motion sequences. We utilize motion imputing on the interaction model to achieve a smooth transition. In particular, we impute the start of the interaction sequence from the entire sequence obtained by the (guided) grasping model. Intuitively, this adjusts the generative process for interaction stage based on the observed grasping motion. To perform the imputation, we define the projection that resizes the hand poses from the grasping stage to hand-object poses from the interaction stage by filling in zeros, both temporally and spatially. Similarly, represents the imputation regions of the grasp sequence on . At each denoising step, the imputed denoised output is given by
where is elementwise multiplication. Our proposed technique, dubbed subsequence imputing, leads to smoother transition between grasp interaction compared to single frame transition (see Sec. 5.6).
@@ 4.3. Grasp Guidance
To add controllability to the generation and improve generalization to unseen objects, we introduce grasp guidance for the grasping stage, which can be optionally added at inference time via the goal function (see Sec. 3 and Eq. 2). We leverage a single grasp frame as target grasp (hand pose) for the last frame of the grasping stage. It provides a prior about how and with which hand an object should be grasped. Such a grasp can be obtained from grasp generation (Zhang et al., 2024a) or from an off-the-shelf hand-object pose estimator (Pavlakos et al., 2024) as demonstrated in Sec. 5.7. We aim to minimize the goal function and approximate the gradient of the guidance signal in Eq. 2 as
where is the hand(s) pose in the last frame of the grasping sequence after denoising and the cost term minimizes the distance of the hand(s) pose to the target grasp. Depending on the availability, the guidance signal may include one or two hands. Since the diffusion process is a denoising model of a motion sequence, it relates individual frames to other frames in the sequence. Hence, by backpropagating through the diffusion model (via autodiff), we can compute a dense guidance signal for all frames, even if only sparse guidance signals (one in our case) are provided.
4.4. Detailed Text Descriptions
@@
Our model generates hand-object interactions based on textual inputs. To the best of our knowledge, there is currently no HOI dataset that contains detailed textual descriptions. For instance, the GRAB dataset (Taheri et al., 2020) only provides categorical action labels. To address this, we auto-generated sentences using the template “the person ¡verb¿ + ¡object¿” (Ghosh et al., 2023) (referred to as “simple” text). However, these sentences lack detailed information, which limits the controllability of the model via textual inputs. Therefore, we contribute carefully annotated textual descriptions of GRAB (referred to as “detailed” sentences). We instructed our annotators to carefully watch each ground truth video and describe the actions in three distinct stages of pre-action, action, and post-action. These descriptions are further augmented with hand and positional information. An example from the dataset is as follows: ”The person picks up the apple with the right hand, passes it to their left hand, and places it back with their right hand”. These detailed descriptions enhance the dataset, enabling more accurate generation and precise control of HOI as we show in Sec. 5.8.
| Method | Backbone | SD [] () | OD []() | IV [] () | ID [] () | CR |
| Unseen subject split | IMoS (Ghosh et al., 2023) | CVAE | 0.002 | 0.149 | 7.14 | 11.47 | 0.05 | 0.588 | |
| DiffH2O | Transformer | 0.088 | 0.185 | 6.65 | 8.39 | 0.067 | 0.810 | ||
| DiffH2O | UNet | 0.109 | 0.188 | 6.02 | 7.92 | 0.064 | 0.875 | ||
| Unseen object split | IMoS* (Ghosh et al., 2023) | CVAE | 0.002 | 0.132 | 10.38 | 12.45 | 0.048 | 0.581 | |
| DiffH2O | Transformer | 0.133 | 0.185 | 7.99 | 10.87 | 0.073 | 0.803 | ||
| DiffH2O | UNet | 0.134 | 0.179 | 9.03 | 11.39 | 0.086 | 0.837 |
@@ 5.1. Baselines
Our method focuses on synthesizing detailed hand-object motion based on text input. To the best of our knowledge, DiffH2O is the first method to tackle this problem. While there is no direct baseline for hand-object motion synthesis based on language, the closest to our approach is IMoS (Ghosh et al., 2023), which focuses on text-based whole-body human-object interaction synthesis. IMoS first generates a body articulation based on an instruction label (action + object), and then optimizes for the object pose using a grasp heuristic. Since our paper focuses on hand articulations and object motions, we omit the full-body movements from our evaluations.
To further compare DiffH2O against human-body diffusion models, such as (Tevet et al., 2023) and (Karunratanakul et al., 2023), we adopted them to the HOI setting. All variants use our proposed pose representation. MDM (Tevet et al., 2023) does not use guidance, and for GMD (Karunratanakul et al., 2023), we use gradients to guide the model towards the grasp frame, but omit the 2D-trajectory generation stage because it is specific to human motions.
@@ 5.2. Data
5.2.1. GRAB
We utilize the subject-based split of the GRAB dataset (Taheri et al., 2020), which contains 1335 sequences, proposed in IMoS to run a direct comparison. However, as this split does not contain unseen objects, we also create a new split featuring unseen objects based on object similarity and semantics (see supplemental material). To assess our model’s ability to generalize to unseen classes, we introduce a new split comprising unseen objects that were not present in training. Our train and test split contains 1125 and 210 sequences, respectively, which we mirror to a total of 2250 and 420 sequences. To split the GRAB sequences into the grasping and interaction stages for training, we apply the heuristics described in Section 4.2.1 Our test set is composed of the following objects: “apple, mug, train, elephant, alarm clock, small pyramid, medium cylinder and large torus”. Note that the training set includes pyramids, cylinders, and torus in other sizes. We exclude small pyramids, medium cylinders, and large torus from the training set to test our models’ ability to generalize to different sizes of the same objects To evaluate grasp guidance in a controlled manner, we utilize the frame of the unseen test split that indicates the transition between the grasp and interaction stage. Note that the grasp reference contains both hands in this case, even if one hand is not grasping the object.
5.2.2. HO-3D
To further evaluate guidance and object generalization, we estimate poses on the HO-3D dataset (Hampali et al., 2020) and guide our model to an estimated target hand pose.
@@ 5.3. Evaluation Metrics
We report several physics-based metrics and motion diversity metrics following previous works (Ghosh et al., 2023; Tendulkar et al., 2023; Taheri et al., 2022; Braun et al., 2024). Please see supplementary material for additional metrics based on action features.
5.3.1. Physics Based Metrics
@@
We report the interpenetration volume (IV) as the number of MANO vertices penetrating the object mesh and the maximum interpenetration depth (ID). Additionally, we compute the contact ratio (CR) as the average ratio of hand vertices within 5mm of the object mesh.
5.3.2. Motion Metrics
@@
We report the sample diversity (SD) between wrist trajectories in motion space. We sample the same input conditions 5 times, compute the pairwise Euclidean distance between the samples and report the mean SD over all test prompts. We also provide the overall diversity (OD), which measures the pairwise Euclidean distance between all test samples. Furthermore, we report the action recognition accuracy (AR) as in (Ghosh et al., 2023).
5.3.3. Guidance Metrics
@@
To assess how well the models adhere to the guidance signal, we report the mean error between the predictions and the reference grasp (GE), and the ratio at which the diffusion model maintains the correct handedness according to the reference (HA). Furthermore, we evaluate the wrist velocities at the transition between grasping and interaction phase ().
@@
5.4. Comparison to the State-of-the-Art
We conduct a set of experiments to compare to relevant baselines quantitatively. We present qualitative examples of our method in Fig. 4 and failure cases in Fig. 5.
5.4.1. Setting 1 - Interaction Only
We compare our approach against the closest existing work, IMoS, in their interaction-only setting and report results in Tab. 1. To ensure a fair comparison with IMoS, we evaluate our model without the grasp guidance and without our detailed text annotations. We test DiffH2O with two backbones, transformer as in (Tevet et al., 2023) and UNet as in (Karunratanakul et al., 2023). Our model significantly outperforms IMoS across all physics and motion diversity metrics. In particular, our method yields motions that are significantly more diverse (higher SD and OD), exhibit less interpenetration (lower IV and ID), and align better with the action type (higher AR). Note that IMoS assumes the availability of a ground-truth test motion to initialize the first few frames, while ours predicts the entire sequence. Yet, our approach achieves considerable performance improvements. While IMoS provides compelling realistic full body synthesis results, fine-grained finger motions and object motions lag behind the realism and diversity demonstrated by our method, as shown in Fig. 3 and in our supplementary video. In Section 5.5, we further provide a perceptual study to compare our method against the baseline.
5.4.2. Setting 2 - Grasping + Interaction
In this experiment, we focus on generating entire sequences of grasping and interaction, evaluating how well the models can produce high-quality motions without physical artifacts and adhere to guidance signals. Results in Tab. 2 show our method outperforms human motion diffusion baselines (MDM/GMD) with better grasp accuracy (GE), handedness (HA), and naturalness (IV). This validates our temporal two-stage design with grasp guidance and subsequence imputing over previous diffusion models for the setting of hand-object interactions.
5.5. Perceptual User Study
To evaluate the visual quality of our motions, we conducted a perceptual user study on a set of 21 participants, who were randomly assigned one of four sets, each consisting of 30 randomly selected sequences. We compared our approach against IMoS. We displayed results from our method and IMoS side-by-side in random order, along with the input sentence description used to synthesize motions. We asked the following questions to the participants: “Which sequence is more realistic?” and “Which sequence has more more variety in motion?”. We defined a sequence to be more realistic if the overall motion looks human-like, e.g., if the object is grasped realistically and there are less artifacts such as floating objects or interpenetration. Moreover, a sequence has more variety if the object is manipulated multiple times and in distinct ways. In 63.1 of the responses, participants selected our method as the most realistic compared to IMoS. Even more distinctively, in 72.9 of the responses, participants favored our method to contain more variety than IMoS. We obtained statistically significant Fleiss’ kappa scores (Fleiss, 1971) of 0.34 and 0.43 for Realism and Diversity, respectively, at p¡0.05. In our investigation, our model showed low realism but high diversity for the “pass” action, influenced by users perceiving “pass” action label in the GRAB dataset as transferring an object between hands, not to another person. In the “offhand” action category of GRAB, where an object is passed from one hand to another (see Figure 3 and video 02:51), our model excelled in realism, while IMoS exhibited artifacts. This is due to the fact that IMoS determines the hand which is in contact with the object according to ground-truth, which leads to a jump of the object for the “off-hand” category.